# Applying a STEM Ways of Thinking Framework for Student-generated Engineering Design-based Physics Problems

### v1

@@ 5.3 Science-Based Thinking (SBT)
In our rubric (see Appendix D), we asked students to engage in what we referred to as science thinking [9], which included the following components: (i) the use of at least two of the three fundamental principles—momentum, energy, and angular momentum—along with a justification for their relevance; (ii) identifying the assumptions and approximations made; (iv) identifying the system and surroundings specific to the problem (systems thinking); and (v) employing modeling by explicitly stating and justifying the use of a model (e.g., point-particle) in the solution.
@@
In this study, we distinguish between physics principles and physics concepts, although some overlap is inevitable. By physics principles, we specifically refer to the three fundamental principles: momentum, energy, and angular momentum. In contrast, physics concepts include terms such as force, air resistance, and buoyancy, among others. We find it useful to adopt this distinction as our course textbook [47] follows a principle-based approach. Some might view this as a bookkeeping measure. Although our approach was informed by our recent study [9], the reader will notice a significant increase in the level of detail in the coding schema for this study.
@@
While only a couple of teams addressed the notion of ‘system and surroundings’ within their problem statements, most teams addressed this while progressing through their solutions. Team-12, in the course of writing down the related equations, explicitly stated:
@@
Group discussions across teams tended to be more general and, at times, reiterated points that were documented in the reports. However, the value of these discussions should not be overlooked, as they provide insight into how ideas evolved within teams and highlight the interests and contributions of different members. In contrast, the written reports, while lacking the dynamic nature of discussions, exhibited a higher level of detail, reflecting a more structured engagement with the design process. For example, Team-01, during group discussions, reflected on their iterations:
“And then from there we iterated through multiple degrees for the first part and then a multitude of impact locations for the second part and graphed our findings”, adding “the confidence in our analysis spreads to two primary aspects of baseball. The optimum angle and the point of impact on the ball. By understanding these factors, we’re going to improve the performance and make more informed decisions and adjust our techniques and yield better results”.
Evidently, the team is trying to convey that the range of the projectile depends on both the angle of projection relative to the horizontal and the point where the bat impacts the ball, and they outline how they intend to calculate the projectile’s range. Although the conversation snippet may seem a bit unclear, the team presented calculations to support their reasoning in their reports by employing the momentum update principle.
We expected students to apply at least two of the three fundamental principles but allowed flexibility to accommodate their natural approaches to applying physics concepts. The majority of teams applied either the momentum principle, the energy principle, or both, while fewer teams incorporated the angular momentum principle (see Figure 20 in Appendix B). This may be attributed to the nature of the problems chosen by teams or the fact that angular momentum is introduced later in the course (around week 10), leaving students with less time to assimilate and apply related ideas. Team-07 outlined their use of momentum principle incorporating the use of mathematics as follows:
@@
In progression toward their solution, in their reports, they generated a graph and table adding a level of detail to their discussions. Team-11 explicitly articulated their assumption (negligible air resistance) and approximation (the cord obeys Hooke’s law), along with their initial conditions:
 “We assume air resistance is negligible and the cord obeys Hooke’s law, and that the person will start and end at rest”. They further clarified their system definition, stating: “The system is the person, cord, and the Earth, and we have nothing in the surroundings because no work is done on the system”.
@@
One notable concern was the lack of attention to numerical precision. None of the teams presented numerical results with appropriate significant figures, likely due to an over-reliance on Python for automated calculations. Additionally, since students were not creating physical models or working with actual data, they may not have perceived significant figures as particularly relevant. This combination of factors likely contributed to the oversight in numerical precision. Additionally, a significant number of teams neglected to include correct units when reporting results—an essential expectation in scientific communication. The numerical results presented by Team-02 (see Figure 14) exemplify this trend, which was prevalent across almost all project reports (see Figure 20 in Appendix B).
@@
While they calculated the moment of inertia, they did not specify the axis around which the calculation was made, which is an important detail for accuracy in physical modeling. It is possible it was another instance oversight. Similarly, Team-14, when analyzing the forces on the yo-yo, stated:
@@
This suggests a misconception about the forces at play at the lowest point of the yo-yo’s motion. In reality, the net force is not zero at the lowermost point. It appears the students may have thought momentary rest may imply zero acceleration. For a more accurate understanding of the yo-yo’s behavior, we refer the reader to the insightful article by Minkin and Sikes, which explains the phenomenon of how a yo-yo sleeps at the lowest point of its trajectory [137]. This example underscores the complexity of real-world problems, demonstrating that even seemingly straightforward scenarios can invite deeper investigation and further research [138]. Despite this conceptual oversight, Team-14 did demonstrate a correct understanding in another aspect, stating that “the yo-yo won’t slip on the string”, which is a key insight to use the rolling constraint equation $v_{\text{cm}}=r\omega$. The importance of deep understanding of physics concepts is clear in these examples. Without careful attention to accuracy and clarity, even well-intentioned design efforts can falter, making it essential for students to avoid conceptual flaws and ensure their scientific understanding underpins their design work.
@@
Given that this is a physics course, a systematic approach to iterations could include requiring students to explain whether each iteration meant an incremental shift in the application of physics concepts, thereby leading to a batter solution.
@@ 5.4 Mathematics-Based Thinking (MBT)
For MBT, we analyzed students’ use of algebra, including basic trigonometry and geometry, in their calculations. Despite this being a calculus-based course, only one team (Team-04) explicitly applied calculus in their analysis. This team formulated coupled first-order differential equations in two dimensions to account for air friction in projectile motion, implemented a Python script to solve the equations, and graphically represented the results (see Figure 17). Notably, the teams’ Python implementation and mathematical approach extended well beyond the curricular expectations, illustrating how students can sometimes exceed anticipated learning boundaries. Additionally, a few teams employed an iterative approach to updating position and momentum, implicitly engaging with principles of differential calculus (see Figure 16).
A few teams reasoned qualitatively during discussions and provided detailed calculations in their reports. Team-04 articulated their reasoning behind using differential equations as follows:
@@
The reader may notice how students’ language weaves together both physics and mathematics. Notably, this team did not explicitly specify the axis when referring to the moment of inertia, either in their discussions or reports. This example also reinforces our choice of data sources —group discussions and reports. Our experience analyzing group discussions supports the value of taking a more dialogic approach to promote student learning, ultimately fostering a “culture of talk” [142] in our classrooms. Just as writing supports Writing to Learn (WTL) [143], our findings suggest that discussions have the potential promote what we call Talking to Learn (TTL). Further research is needed to develop scaffolds and frameworks that foster Hammer’s notion of “productive” [17] discussions, or “rich talk” as put forth by City [144, p.13].
@@
A key gap in students’ MBT was the limited application of calculus in their problem-solving, despite its relevance to the course. While some teams implicitly used differential calculus concepts through iterative position and momentum updates, there was little explicit engagement with calculus-based reasoning. Additionally, Python was primarily used for graphing, with very few teams employing it to generate tables, suggesting a narrow scope of computational tool use.
To address these gaps, targeted instructional strategies can be implemented in future iterations. Inclusion of calculus-based tasks in the course, or in laboratory tasks if possible, could help students recognize its relevance and applicability. Similarly, incorporating assignments that require students to use Python for both graphing and tabular data representation may encourage broader engagement with computational tools. To strengthen proportional reasoning, guided reflection activities or structured peer discussions could help students articulate and apply mathematical relationships more effectively. Finally, embedding opportunities for students to collect, analyze, and integrate experimental or database-sourced data into their projects could foster deeper engagement with real-world problem-solving and reinforce the role of empirical data in mathematical modeling.
In future iterations of our ED-based project, to further support MBT, we see scope for integrating elements of the ACER framework proposed by Wilcox et al. [145, p.4]Activation—problem cues trigger relevant mathematical tools; Construction—students develop simplified mathematical models using equations and representations; Execution—students manipulate mathematical expressions while considering their physical meaning; Reflection—students evaluate their results, checking for errors and assessing physical insights.
@@ 5.5 Metacognitive Reflection (MER)
While reflective statements appeared in both transcripts (see Figure 3) and reports, our focus on future pedagogical iterations led us to identify the most relevant data segment for analysis. In the conclusion section of their final reports, we asked students:
@@
Student responses to this prompt provided valuable insights, as written reports represented collective reasoning of the team rather than individual contributions in discussions. Our goal was to assess whether this prompt elicited meaningful reflections that could inform future course design. For our analysis we used the written reports of all the 14 teams. See Table 4.
@@
In summary, students’ responses to this question provided little new insight into the teams’ thinking, beyond what had already been stated by the teams. It is possible that they were fatigued by the end of the report or that the wording of the prompt did not effectively encourage deeper reflection.
@@
A more structured approach to fostering metacognitive reflection may be necessary, as simply asking students to reflect on their solutions does not appear to be effective. As Hacker puts it, students “should be directed toward specific, higher-level analyses” [147, p.79]. MER has the potential to reveal the interconnectedness of students’ Ways of Thinking. While interconnectedness of the ways of thinking is not intentionally explored in this paper, it will be the focus of our third paper. For the next iteration, we plan to use prompts specifically designed to elicit this interconnectedness, encouraging students to critically assess their solutions by considering physics, math, coding, assumptions, feasibility, and innovation. This aligns with our goal of fostering deeper metacognitive engagement and justification of design decisions. We also see scope for incorporating ideas from Science Writing Heuristic (SWH) [143, p.5], which encourages reflective writing at the end of a lab report.
@@ 5.6 Computational Thinking (CT)
Second, we applied a deductive coding approach, guided by the CT framework outlined by Shute et al. [71], which provided a structured lens through which to interpret specific elements of CT. This deductive layer enabled a systematic assessment of how closely students’ work aligned with established CT principles, facilitating comparison with standard CT benchmarks. A coding sheet, similar to that in Figure 10, was used in the early stages but is not reproduced here. Table LABEL:tab:deductive_CT_coding in Appendix C produces a summary for the entire data set.
@@
While standard Python libraries, such as numpy (used for efficient numerical calculations of vectors and matrices), pandas (used for data manipulation and analysis), and matplotlib (used for data visualization in graphs and charts), had been introduced in prior labs, only a few groups utilized numpy and pandas, whereas matplotlib (see Figure 17) saw a relatively higher adoption rate. Notably, only one team applied an advanced library like scipy (used for advanced computations involving differentiation and integration), suggesting that deeper CT practices were less common. These patterns indicate that, while students demonstrated a baseline familiarity with Python’s tools for data handling and visualization, advanced computational approaches were relatively rare.
@@
Data visualization through graphs (see Figure 17) and tables was moderately represented, with several groups recognizing the value of visually presenting their data or findings, although the sophistication of these visualizations varied. Most teams used built-in basic commands, such as print (see Figures 14 and 15), for either displaying the output, or perhaps as a debugging tool. The incorporation of physics formulas was nearly universal, yet varied in complexity, with some groups embedding straightforward calculations. Notably, some groups had typed in incorrect formulas, probably due to oversight.
As mentioned in Section (ii), due to a possible over-reliance on Python coding, teams paid little attention to significant digits when reporting numerical results—sometimes extending to more than 10 digits! (see Figure 14). Without controlling for precision, students may miss important learning opportunities related to numerical precision, an important aspect in science learning. This example also highlights how science-based thinking and computational thinking are interconnected.

### v2

@@ 5.3 Science-Based Thinking (SBT)
In our rubric (see Appendix D), we asked students to engage in what we referred to as science thinking [9], which included the following components: (i) the use of at least two of the three fundamental principles—momentum, energy, and angular momentum—along with a justification for their relevance; (ii) identifying the assumptions and approximations made; (iv) identifying the system and surroundings specific to the problem (systems thinking); and (v) employing modeling by explicitly stating and justifying the use of a ‘model’ (e.g., point-particle) in the solution. Since teams did not detail upon the choice of their model directly, for the purposes of this study, we found it suitable and sufficient to merge point (v) into (iv) under ‘systems and surroundings’. Furthermore, points (iv) and (v) involve deeper notions such as ‘systems thinking’ [140, 141, 142] and ‘modeling’ in science [143, 144], the discussion of which is well beyond our immediate pedagogical goals.
@@
In this study, we distinguish between physics principles and physics concepts, although some overlap is inevitable. By physics principles, we specifically refer to the three fundamental principles: momentum, energy, and angular momentum that are presented in the textbook [47] used in this course. In contrast, physics concepts include terms such as force, air resistance, and buoyancy, among others. We find it useful to adopt this distinction as our course textbook [47] follows a principle-based approach. Some might view this as a bookkeeping measure. Although our approach was informed by our recent study [9], the reader will notice a significant increase in the level of detail in the coding schema for this study.
@@
While only a couple of teams addressed the notion of ‘system and surroundings’ within their problem statements, a majority of teams addressed this while progressing through their solutions. Team-12, in the course of writing down the related equations, explicitly stated:
@@
Group discussions across teams tended to be more general and, at times, reiterated points that were documented in the reports. However, the value of these discussions should not be overlooked, as they provide insight into how ideas evolved within teams and highlight the interests and contributions of different members – an aspect which may be of interest for some researchers. In contrast, the written reports, while lacking the dynamic nature of discussions, exhibited a higher level of detail, reflecting a more structured engagement with the design process. For example, Team-01, during group discussions, reflected on their iterations:
“And then from there we iterated through multiple degrees for the first part and then a multitude of impact locations for the second part and graphed our findings”, further adding “the confidence in our analysis spreads to two primary aspects of baseball. The optimum angle and the point of impact on the ball. By understanding these factors, we’re going to improve the performance and make more informed decisions and adjust our techniques and yield better results”.
Evidently, the team is trying to convey that the range of the projectile depends on both the angle of projection relative to the horizontal and the point where the bat impacts the ball, and they outline how they intend to calculate the projectile’s range. Although the conversation snippet may seem a bit unclear, the team presented calculations to support their reasoning in their reports by employing the momentum update principle employing Python code (for the code, see Figure 16).
We expected students to apply at least two of the three fundamental principles but allowed flexibility to accommodate their natural approaches to applying physics concepts. A majority of teams applied the momentum principle and/or the energy principle, while fewer teams incorporated the angular momentum principle (see Figure 20 in Appendix B). This may be attributed to the nature of the problems chosen by teams or the fact that angular momentum is introduced later in the course (around week 10), leaving students with less time to assimilate and apply related ideas. Team-07 outlined their use of momentum principle incorporating the use of mathematics as follows:
@@
In progression toward their solution, in their reports, they generated a graph and table adding a level of detail to their discussions.
Team-11 explicitly articulated their assumption (negligible air resistance) and approximation (the cord obeys Hooke’s law), along with their initial conditions:
 “We assume air resistance is negligible and the cord obeys Hooke’s law, and that the person will start and end at rest”. They further clarified their system and surroundings, stating: “The system is the person, cord, and the Earth, and we have nothing in the surroundings because no work is done on the system”.
@@
One notable concern was the lack of attention to numerical precision. None of the teams presented numerical results with appropriate significant figures, likely due to an over-reliance on Python for automated calculations. Additionally, since students were not creating physical models or working with data from direct experimentation, they may not have perceived significant figures as particularly relevant. This combination of factors likely contributed to the oversight in numerical precision. Additionally, a significant number of teams neglected to include correct units when reporting results—an essential expectation in scientific communication. The numerical results presented by Team-02 (see Figure 14) exemplify this trend, which was prevalent across almost all project reports (see Figure 20 in Appendix B).
@@
While they calculated the moment of inertia, they did not specify the axis around which the calculation was made, which is an important detail for accuracy in physical modeling. It is possible it was another instance of oversight. Similarly, Team-14, when analyzing the forces on the yo-yo, stated:
@@
This suggests the conceptual resources [145] that students are activating to describe the forces at play at the lowest point of the yo-yo’s motion. In reality, the net force is not zero at the lowermost point. It appears the students may be relating momentary rest as implying zero acceleration. For a more accurate understanding of the yo-yo’s behavior, we refer the reader to the insightful article by Minkin and Sikes, which explains the phenomenon of how a yo-yo sleeps at the lowest point of its trajectory [146]. This example underscores the complexity of real-world problems, demonstrating that even seemingly straightforward scenarios can invite deeper investigation and further research [147]. Despite this conceptual oversight, Team-14 did demonstrate a correct understanding in another aspect, stating that “the yo-yo won’t slip on the string”, which is a key insight to use the rolling constraint equation $v_{\text{cm}}=r\omega$.
In our previous study [9], we advocated for strengthening the design-science connection. A strong understanding of science concepts can enhance design, while weak scientific reasoning can derail it. Engaging in a critical evaluation of results may help mitigate such issues. For instance, Team-05 determined a spring constant (about 3500 N/m) which they thought was unrealistically high stating:
“roughly 3 - 4 times more than a heavy bow shot would be pulled”,
but demonstrated thoughtfulness by referring to online sources and acknowledging the discrepancy. In contrast, Team-09, despite having a well-conceived design, used incorrect formulas, leading to misleading results that they did not verify. They also assumed a spring constant of 100 N/m, which seemed excessively high for a toy. Similarly, Team-12, though correctly applying physics concepts, made a calculation error that resulted in at least two issues: (1) a formula that was dimensionally incorrect, which they failed to notice, and (2) a numerical result for the skater’s speed (about 11 m/s) that exceeded that of an Olympic sprinter! Not to be left behind (pun intended), Team-02 calculated their golf cart’s speed to be around 9 m/s but did not critically evaluate the plausibility of such a high speed. To be fair to the students, it is possible that they did not notice these issues due to lack of sufficient time to review their calculations and reflect on their results.
The importance of deep understanding of physics concepts is clear in these examples. Without careful attention to accuracy and clarity, even well-intentioned design efforts can falter, making it essential for students to avoid conceptual flaws and ensure their scientific understanding underpins their design work.
@@
Given that this is a physics course, a systematic approach to iterations could include requiring students to explain whether each iteration meant an incremental shift in the application of physics concepts, thereby leading to a better solution.
@@ 5.4 Mathematics-Based Thinking (MBT)
Since teams used Python for calculations, readers may notice an overlap with computational thinking, which we address separately in Section 5.6. However, this overlap will not hinder our discussion, as our focus here remains on mathematical thinking.
@@
For MBT, we analyzed students’ use of algebra, including basic trigonometry and geometry, in their calculations. Despite this being a calculus-based course, only one team (Team-04) explicitly applied calculus in their analysis. This team formulated coupled first-order differential equations in two dimensions to account for air friction in projectile motion, implemented a Python script to solve the equations, and graphically represented the results (see Figure 17). Notably, this teams’ Python implementation and mathematical approach extended well beyond the curricular expectations, illustrating how students can sometimes exceed anticipated learning boundaries. A smaller fraction of teams employed an iterative approach to updating position and momentum. While the calculations may initially appear algebraic, they implicitly engage with principles of differential calculus (see Figure 16).
While there were instances of qualitative reasoning during discussions in about half of the teams, only a smaller fraction followed up with additional, detailed calculations in their reports. Team-04 articulated their reasoning behind using differential equations as follows:
@@
The reader may notice how students’ language weaves together both physics and mathematics. Notably, this team did not explicitly specify the axis when referring to the moment of inertia, either in their discussions or reports. This example further reinforces our choice of data sources —group discussions and reports. Our experience analyzing group discussions supports the value of taking a more dialogic approach to promote student learning, ultimately fostering a “culture of talk” [151] in our classrooms. Just as writing supports Writing to Learn (WTL) [152], our findings suggest that discussions have the potential promote what we call Talking to Learn (TTL). Further research is needed to develop scaffolds and frameworks that foster Hammer’s notion of “productive” [17] discussions, or “rich talk” as put forth by City [153, p.13].
@@
A key gap in students’ MBT was the limited application of calculus in their problem-solving, despite its relevance to the course. While some teams implicitly used differential calculus concepts through iterative position and momentum updates, there was little explicit engagement with calculus-based reasoning. Python was primarily used for graphing, with very few teams employing it to generate tables.
To address these gaps, targeted instructional strategies can be implemented in future iterations. Inclusion of calculus-based tasks in the course, or in laboratory tasks if possible, could help students recognize its relevance and applicability. Incorporating assignments that require students to use Python for both graphing and tabular data representation could encourage broader engagement with computational tools. Additionally, students should be encouraged to analyze graphs and tables more deeply, identifying patterns and discussing the underlying science to better inform their design decisions.or patterns, and discussing the underlying science so as to inform their design decisions.
To strengthen proportional reasoning, guided reflection activities or structured peer discussions could help students articulate and apply mathematical relationships more effectively. Finally, embedding opportunities for students to collect, analyze, and integrate experimental or database-sourced data into their projects could foster deeper engagement with real-world problem-solving and reinforce the role of empirical data in mathematical modeling.
In future iterations of our ED-based project, to further support MBT, we see scope for integrating elements of the ACER framework proposed by Wilcox et al. [154, p.4]Activation—problem cues trigger relevant mathematical tools; Construction—students develop simplified mathematical models using equations and representations; Execution—students manipulate mathematical expressions while considering their physical meaning; Reflection—students evaluate their results, checking for errors and assessing physical insights.
@@ 5.5 Metacognitive Reflection (MER)
While reflective statements appeared in both transcripts (see Figure 3) and reports, our focus on future pedagogical iterations led us to identify the most relevant portion of data for analysis. In the conclusion section of their final reports, we asked students:
@@
Student responses to this prompt provided valuable insights, as written reports represented collective reasoning of the team rather than individual contributions in discussions. Our goal was to assess whether this prompt elicited meaningful reflections that could inform future course design. Hence, we used teams’ responses to the above prompt from the written reports of all 14 teams. See Table 4.
@@
In summary, students’ responses to this question provided little new insight into the teams’ thinking, beyond what had already been stated by the teams. It is possible that they were fatigued by the end of the report [156] or that the wording of the prompt did not effectively encourage deeper reflection.
@@
A more structured approach to fostering metacognitive reflection may be necessary, as simply asking students to reflect on their solutions does not appear to be effective. As Hacker puts it, students “should be directed toward specific, higher-level analyses” [157, p.79]. MER has the potential to reveal the interconnectedness of students’ Ways of Thinking. While interconnectedness of the ways of thinking is not intentionally explored in this paper, it will be the focus of our third paper. For the next iteration, we plan to use prompts specifically designed to elicit this interconnectedness, encouraging students to critically assess their solutions by considering physics, math, coding, assumptions, feasibility, and innovation. This aligns with our goal of fostering deeper metacognitive engagement and justification of design decisions. We also see scope for incorporating ideas from Science Writing Heuristic (SWH) [152, p.5], which encourages reflective writing at the end of a lab report.
@@ 5.6 Computational Thinking (CT)
Second, we applied a deductive coding approach, guided by the CT framework outlined by Shute et al. [71], which provided a structured lens through which to interpret specific elements of CT. This deductive layer enabled a systematic assessment of how closely students’ work aligned with established CT principles, facilitating comparison with standard CT benchmarks. A coding sheet, similar to that in Figure 10, was used in the early stages but is not reproduced here. Table LABEL:tab:deductive_CT_coding in Appendix C produces a summary of findings for the entire data set.
@@
While standard Python libraries, such as numpy (used for efficient numerical calculations of vectors and matrices), pandas (used for data manipulation and analysis), and matplotlib (used for data visualization in graphs and charts), had been introduced in prior labs, only a few teams utilized numpy and pandas, whereas matplotlib (see Figure 17) saw a relatively higher adoption rate. Notably, only one team applied an advanced library like scipy (used for advanced computations involving differentiation and integration), suggesting that deeper CT practices were less common. These patterns indicate that, while students demonstrated a baseline familiarity with Python’s tools for data handling and visualization, advanced computational approaches were relatively rare.
@@
Data visualization through graphs (see Figure 17) and tables was moderately represented, with several teams recognizing the value of visually presenting their data or findings, although the sophistication of these visualizations varied. Most teams used built-in basic commands, such as print (see Figures 14 and 15), for either displaying the output, or perhaps as a debugging tool. The incorporation of physics formulas was nearly universal, yet varied in complexity, with some teams embedding straightforward calculations. Notably, some teams had typed in incorrect formulas, probably due to oversight.
As mentioned in Section (ii), due to a possible over-reliance on Python coding, teams paid little attention to significant digits when reporting numerical results—sometimes extending to 17 digits! (see Figure 14). Without controlling for precision, students may miss important learning opportunities related to numerical precision, an important aspect in science learning. This example also highlights how science-based thinking and computational thinking are interconnected.